OpenStudy (anonymous):

Integrate ∫ (cos2x-cos2α)/ (cosx-cosα) dx, where α is independent of x.

OpenStudy (blockcolder):

\[\cos A-\cos B=-2\sin\left(\frac{A+B}{2}\right)\sin\left(\frac{A-B}{2}\right)\] Use this on the numerator and denominator.